Marge is twice as old as consuelo.the sum of their ages seven years ago was 13 how old are they now ?
Answer by algebrahouse.com(1659)

"Marge is twice as old as consuelo.the sum of their ages seven years ago was 13 how old are they now ?

x = Consuelo's age now
2x = Marge's age now

x - 7 = Consuelo's age 7 yrs ago
2x - 7 = Marge's age 7 yrs ago

x - 7 + 2x - 7 = 13 {sum of ages 7 yrs ago was 13}
3x - 14 = 13 {combined like terms}
3x = 27 {added 14 to both sides}
x = 9 {divided both sides by 3}
2x = 18 {substituted 9, in for x, into 2x}

Consuelo is 9 now
Marge is 18 now
www.algebrahouse.com